# On Gottschalk’s surjunctivity conjecture for non-uniform cellular automata

### v1

@@ Abstract
Gottschalk’s surjunctivity conjecture for a group $G$ states that it is impossible for cellular automata (CA) over the universe $G$ with finite alphabet to produce strict embeddings of the full shift into itself. A group universe $G$ satisfying Gottschalk’s surjunctivity conjecture is called a surjunctive group. The surjunctivity theorem of Gromov and Weiss shows that every sofic group is surjunctive. In this paper, we study the surjunctivity of local perturbations of CA and more generally of non-uniform cellular automata (NUCA) with finite memory and uniformly bounded singularity over surjunctive group universes. In particular, we show that such a NUCA must be invertible whenever it is reversible. We also obtain similar results which extend to the class of NUCA a certain dual-surjunctivity theorem of Capobianco, Kari, and Taati for CA.
Keywords:
Gottschalk’s conjecture surjunctivity sofic groupcellular automata non-uniform cellular automata asynchronous cellular automata reversibility
@@ 1 Introduction
We generalize our results to a certain class of global perturbations of CA with uniformly bounded singularity. Essentially by definition and basic properties of CA, every reversible CA with finite memory over a surjunctive group universe must be invertible (see Section 1.1 for the exact definitions). Gottschalk’s surjunctivity conjecture amounts to say that every group is surjunctive. Hence, our results extend Gottschalk’s surjunctivity conjecture by showing that over surjunctive group universes, reversible NUCA with finite memory and uniformly bounded singularity must be invertible (see Theorem A and Theorem C).
@@ 1.1 Basic definitions
We recall notions of symbolic dynamics. Given a discrete set $A$ and a group $G$, a configuration $x\in A^{G}$ is a map $x\colon G\to A$. Two configurations $x,y\in A^{G}$ are asymptotic if $x|_{G\setminus E}=y|_{G\setminus E}$ for some finite subset $E\subset G$. The Bernoulli shift action $G\times A^{G}\to A^{G}$ is defined by $(g,x)\mapsto gx$, where $(gx)(h)=x(g^{-1}h)$ for $g,h\in G$, $x\in A^{G}$. The full shift $A^{G}$ is equipped with the prodiscrete topology. For each $x\in A^{G}$, we define $\Sigma(x)=\overline{\{gx\colon g\in G\}}\subset A^{G}$ as the smallest closed subshift containing $x$.
Following von Neumann and Ulam [24], we define a CA over the group $G$ (the universe) and the set $A$ (the alphabet) as a $G$-equivariant and uniformly continuous self-map $A^{G}\righttoleftarrow$ [7], [17]. One usually refers to each element $g\in G$ as a cell of the universe. Then every CA is uniform in the sense that all the cells follow the same local transition map. More generally, when different cells can evolve according to different local transition maps to break down the above uniformity, we obtain NUCA [11], [12], [29, Definition 1.1]:
@@
The set $M$ is called a memory and $s\in S^{G}$ the configuration of local transition maps or local defining maps of $\sigma_{s}$. Every CA is thus a NUCA with finite memory and constant configuration of local defining maps. Conversely, we regard NUCA as global perturbations of CA. When $s,t\in S^{G}$ are asymptotic, $\sigma_{s}$ and $\sigma_{t}$ are mutually local perturbations of each other. We also observe that NUCA with finite memory are precisely uniformly continuous selfmaps $A^{G}\righttoleftarrow$. As for CA, such NUCA satisfy the closed image property [29, Theorem 4.4], several decidable and undecidable properties [11], [19], [35] and variants of the Garden of Eden theorem [25], [32].
We will analyze in this paper the relations between the following dynamic properties of $\sigma_{s}$. We say that $\sigma_{s}$ is pre-injective if $\sigma_{s}(x)=\sigma_{s}(y)$ implies $x=y$ whenever $x,y\in A^{G}$ are asymptotic. Similarly, $\sigma_{s}$ is post-surjective if for all $x,y\in A^{G}$ with $y$ asymptotic to $\sigma_{s}(x)$, then $y=\sigma_{s}(z)$ for some $z\in A^{G}$ asymptotic to $x$. By the closed image property, it is known that post-surjectivity implies surjectivity for NUCA with finite memory. We say that $\sigma_{s}$ is stably injective, resp. stably post-surjective, if $\sigma_{p}$ is injective, resp. post-surjective, for every $p\in\Sigma(s)$.
The NUCA $\sigma_{s}$ is said to be reversible or left-invertible if there exists a NUCA with finite memory $\tau\colon A^{G}\to A^{G}$ such that $\tau\circ\sigma_{s}=\operatorname{Id}_{A^{G}}$. A NUCA with finite memory is reversible if and only if it is stably injective (see [29, Theorem A]). We say that $\sigma_{s}$ is invertible if it is bijective and the inverse map $\sigma_{s}^{-1}$ is a NUCA with finite memory [29]. We define also stable invertibility which is in general stronger than invertibility, namely, $\sigma_{s}$ is stably invertible if there exist $N\subset G$ finite and $t\in T^{G}$ where $T=A^{A^{N}}$ such that for every $p\in\Sigma(s)$, we have $\sigma_{p}\circ\sigma_{q}=\sigma_{q}\circ\sigma_{p}=\operatorname{Id}$ for some $q\in\Sigma(t)$. In fact, we will show (see Lemma 1) that every invertible NUCA with finite memory over a finite alphabet and a countable group universe is automatically stably invertible.
@@ 1.2 Main results
The situation for the surjunctivity of NUCA admits some complications. While injective and even reversible NUCA with finite memory may fail to be surjective (see [29, Example 14.3]), results in [29] show that reversible local perturbations of CA with finite memory over an amenable group or an residually finite group universe must be surjective. More generally, we obtain an extension (Theorem A) to cover all reversible, or equivalently stably injective, local perturbations of CA over surjunctive group universes. We also strengthen the conclusion by showing that such NUCA must be stably invertible (see Section 4).
@@
Combining with Theorem 1.1 and the result in [29, Theorem B] where we can replace the stable injectivity by the weaker injectivity condition whenever $G$ is an amenable group, we obtain the following general surjunctivity and invertibility result for NUCA which are local perturbations of CA.
@@
Let $M$ be a finite subset of a countable group $G$. Let $A$ be a finite alphabet and $S=A^{A^{M}}$. Let $s\in S^{G}$ be an asymptotic constant configuration. Then $\sigma_{s}$ is stably invertible in each of the following cases:
@@
Our next results concern a certain dual-surjunctivity version of Gottschalk’s conjecture introduced by Capobianco, Kari, and Taati in [6] which states that every post-surjective CA over a group universe and a finite alphabet is also pre-injective. The authors settled in the same paper [6] the case of sofic group universes. See also [28] for some extensions.
@@
By a similar technique as in the proof of Theorem A, we establish (see Section 5) the following extension of the above result of Capobianco, Kari, and Taati to cover the class of stably post-surjective local perturbations of CA over post-injunctive group universes.
@@
Our main results also hold for global perturbations of CA with uniformly bounded singularity. A NUCA $\sigma_{s}\colon A^{G}\to A^{G}$ has uniformly bounded singularity if for every finite subset $E\subset G$ with $1_{G}\in E=E^{-1}$, we can find a finite subset $F\subset G$ containing $E$ such that the restriction $s|_{FE\setminus F}$ is constant. It is clear that $\sigma_{s}$ has uniformly bounded singularity if $s$ is asymptotically constant. When the universe $G$ is a residually finite group, our notion of uniformly bounded singularity is closely related but not equivalent to the notion of (periodically) bounded singularity for NUCA defined in [29, Definition 10.1]. We obtain the following generalizations of Theorem A and Theorem B (see Section 6 and Section 7).
@@
Let $M$ be a finite subset of a countable surjunctive group $G$. Let $A$ be a finite alphabet and let $S=A^{A^{M}}$. Suppose that $\sigma_{s}\colon A^{G}\to A^{G}$ is stably injective for some $s\in S^{G}$ with uniformly bounded singularity. Then $\sigma_{s}$ is stably invertible.
@@
Let $G$ be a finitely generated post-injunctive group and let $A$ be a finite alphabet. Let $M\subset G$ be finite and let $S=A^{A^{M}}$. Suppose that $\sigma_{s}\colon A^{G}\to A^{G}$ is stably post-surjective for some $s\in S^{G}$ with uniformly bounded singularity. Then $\sigma_{s}$ is stably invertible.
@@ 1.3 Perspectives
Our main results motivate the following natural questions which have an affirmative answer for surjunctive group universes and post-injunctive group universes respectively.
@@
Let $G$ be a group universe and let $A$ be a finite alphabet. Let $M\subset G$ be finite and let $S=A^{A^{M}}$. Suppose that $\sigma_{s}\colon A^{G}\to A^{G}$ is stably injective for some asymptotically constant $s\in S^{G}$. Is $\sigma_{s}$ stably invertible?
@@
Let $G$ be a group universe and let $A$ be a finite alphabet. Let $M\subset G$ be finite and let $S=A^{A^{M}}$. Suppose that $\sigma_{s}\colon A^{G}\to A^{G}$ is stably post-surjective for some asymptotically constant $s\in S^{G}$. Is $\sigma_{s}$ stably invertible?
We can actually show (see Proposition 1 below) that Question 1 and Question 2 are equivalent when restricted to the class of linear NUCA with finite memory. Given a vector space $V$ not necessarily finite, then $V^{G}$ is a vector space with component-wise operations and a NUCA $\tau\colon V^{G}\to V^{G}$ is said to be linear if it is also a linear map of vector spaces or equivalently, if its local transition maps are linear.
@@
Let $G$ be a countable group and let $A$ be a finite vector space alphabet. Question 1 has an affirmative answer for all linear $\mathrm{NUCA}$ $\tau\colon A^{G}\to A^{G}$ with finite memory and uniformly bounded singularity if and only if so does Question 2.
Proof
Let $M\subset G$ be a finite subset and let $s\in S^{G}$ where $S=\mathcal{L}(A^{M},A)$. By linearity, we define $s(g,m)\in\operatorname{End}(A)$ for all $m\in M$ and $g\in G$ by $s(g)(v)=\sum_{m\in M}s(g,m)v(m)$ for all $v\in A^{M}$. By setting $s(g,m)=0$ for $m\in G\setminus M$, we obtain for every $v\in A^{G}$ that $s(g)(v)=\sum_{h\in G}s(g,h)v(h)$. Let $T=\mathcal{L}(A^{*M^{-1}},A^{*})$ where $A^{*}$ is the dual space of $A$. We use the right superscript T to denote the transpose of linear maps. The dual configuration of local defining maps $s^{*}\in T^{G}$ is given by $s^{*}(g,m)\coloneqq s(gm,m^{-1})^{\mathsf{T}}$ for all $g,m\in G$. We define $\sigma_{s}^{*}\coloneqq\sigma_{s^{*}}$ to be the dual linear NUCA of $\sigma_{s}$. It is immediate from the definition that $s$ has uniformly bounded singularity if and only if so does $s^{*}$. By [33, Lemma 5.2], note that $s^{**}=s$ and $\sigma_{s}^{**}=\sigma_{s}$ for all $s\in S^{G}$. We infer from the main results in [33] that
@@
Note that every invertible NUCA with finite memory over a finite alphabet and a countable group universe is automatically stably invertible (see Lemma 1). By identifying $A$ with its dual vector space $A^{*}$, we can now conclude that Question 1 has a positive answer for all linear $\mathrm{NUCA}$ $A^{G}\to A^{G}$ of finite memory with uniformly bounded singularity if and only if so does Question 2. ∎
Linear NUCA with finite memory enjoy similar properties as linear CA such as the shadowing property [2], [4], [20], [3], [21], [5], [27], [26], [8], [33]. The recent result [33, Theorem D.(ii)] states that every stably post-surjective linear NUCA with finite vector space alphabet over a residually finite group universe is invertible whenever it is a local perturbation of a linear CA. When specialized to the class of linear NUCA, Theorem B thus generalizes [33, Theorem D.(ii)] because every residually finite group is sofic and thus post-injunctive by Theorem 1.2. To this end, we postulate another seemingly interesting question.
@@
It is also natural to explore similar questions for the more general class of NUCA with uniformly bounded singularity.

### v2

@@ Abstract
Gottschalk’s surjunctivity conjecture for a group $G$ states that it is impossible for cellular automata (CA) over the universe $G$ with finite alphabet to produce strict embeddings of the full shift into itself. A group universe $G$ satisfying Gottschalk’s surjunctivity conjecture is called a surjunctive group. The surjunctivity theorem of Gromov and Weiss shows that every sofic group is surjunctive. In this paper, we study the surjunctivity of local perturbations of CA and more generally of non-uniform cellular automata (NUCA) with finite memory and uniformly bounded singularity over surjunctive group universes. In particular, we show that such a NUCA must be invertible whenever it is reversible. We also obtain similar results which extend to the class of NUCA a certain dual surjunctivity theorem of Capobianco, Kari, and Taati for CA.
@@ 1 Introduction
We generalize our results to a certain class of global perturbations of CA with uniformly bounded singularity. Essentially by definition and basic properties of CA, every reversible CA with finite memory over a surjunctive group universe must be invertible (see Section 1.1 for the exact definitions). Gottschalk’s surjunctivity conjecture amounts to saying that every group is surjunctive. Hence, our results extend Gottschalk’s surjunctivity conjecture by showing that over surjunctive group universes, reversible NUCA with finite memory and uniformly bounded singularity must be invertible (see Theorem A and Theorem C).
@@ 1.1 Basic definitions
We recall notions of symbolic dynamics. Given a discrete set $A$ and a group $G$, a configuration $x\in A^{G}$ is a map $x\colon G\to A$. Two configurations $x,y\in A^{G}$ are asymptotic if $x|_{G\setminus E}=y|_{G\setminus E}$ for some finite subset $E\subset G$. We say that $x\in A^{G}$ is asymptotically constant if $x$ is asymptotic to some constant configuration $y\in A^{G}$. The Bernoulli shift action $G\times A^{G}\to A^{G}$ is defined by $(g,x)\mapsto gx$, where $(gx)(h)=x(g^{-1}h)$ for $g,h\in G$, $x\in A^{G}$. The full shift $A^{G}$ is equipped with the prodiscrete topology. For each $x\in A^{G}$, we define $\Sigma(x)=\overline{\{gx\colon g\in G\}}\subset A^{G}$ as the smallest closed subshift containing $x$. For every subset $X\subset A^{G}$, we denote the restriction of $X$ to a subset $F\subset G$ by $X_{F}=\{x|_{F}\colon\in X\}$.
Generalizing the first construction of a CA over $\mathbb{Z}^{2}$ by von Neumann [33] (see also [6]), a CA over the group $G$ (the universe) and the set $A$ (the alphabet) is a $G$-equivariant and uniformly continuous self-map $A^{G}\righttoleftarrow$ [10], [26]. One usually refers to each element $g\in G$ as a cell of the universe. Then every CA is uniform in the sense that all the cells follow the same local transition map. More generally, when different cells can evolve according to different local transition maps to break down the above uniformity, we obtain NUCA [16], [17], [38, Definition 1.1]:
@@
The set $M$ is called a memory and $s\in S^{G}$ the configuration of local transition maps or local defining maps of $\sigma_{s}$. Every CA is thus a NUCA with finite memory and constant configuration of local defining maps. Conversely, we regard NUCA as global perturbations of CA. When $s,t\in S^{G}$ are asymptotic, $\sigma_{s}$ and $\sigma_{t}$ are mutually local perturbations of each other. We also observe that NUCA with finite memory as defined in this paper form a proper subset of the set of uniformly continuous selfmaps of the full shift (see also [46][Proposition 14]), e.g. the map $F\colon A^{G}\to A^{G}$ given by $F(x)(g)=x(1_{G})$ for all $x\in A^{G}$ and $g\in G$ is uniformly continuous but it is not a NUCA with finite memory. As for CA, NUCA with finite memory satisfy the closed image property [38, Theorem 4.4], several decidable and undecidable properties [16], [28], [44] and variants of the Garden of Eden theorem [34], [41].
In this paper we analyze the relations between the following dynamic properties of $\sigma_{s}$. We say that $\sigma_{s}$ is pre-injective if $\sigma_{s}(x)=\sigma_{s}(y)$ implies $x=y$ whenever $x,y\in A^{G}$ are asymptotic. Similarly, $\sigma_{s}$ is post-surjective if for all $x,y\in A^{G}$ with $y$ asymptotic to $\sigma_{s}(x)$, then $y=\sigma_{s}(z)$ for some $z\in A^{G}$ asymptotic to $x$. By the closed image property, it is known that post-surjectivity implies surjectivity for NUCA with finite memory. We say that $\sigma_{s}$ is stably injective, resp. stably post-surjective, if $\sigma_{p}$ is injective, resp. post-surjective, for every $p\in\Sigma(s)$.
The NUCA $\sigma_{s}$ is said to be reversible or left-invertible if there exists a NUCA with finite memory $\tau\colon A^{G}\to A^{G}$ such that $\tau\circ\sigma_{s}=\operatorname{Id}_{A^{G}}$. A NUCA with finite memory is reversible if and only if it is stably injective (see [38, Theorem A]). We say that $\sigma_{s}$ is invertible if it is bijective and the inverse map $\sigma_{s}^{-1}$ is a NUCA with finite memory [38]. We define also stable invertibility which is in general stronger than invertibility, namely, $\sigma_{s}$ is stably invertible if there exist $N\subset G$ finite and $t\in T^{G}$ where $T=A^{A^{N}}$ such that for every $p\in\Sigma(s)$, we have $\sigma_{p}\circ\sigma_{q}=\sigma_{q}\circ\sigma_{p}=\operatorname{Id}$ for some $q\in\Sigma(t)$. In fact, we will show (see Lemma 1) that every invertible NUCA with finite memory over a finite alphabet and a countable group universe is automatically stably invertible.
@@
Remark 2
Suppose that the universe $G$ is a countably infinite group and $s\in S^{G}$ is asymptotic to a constant configuration $c\in S^{G}$. By [38][Lemma 8.1] (which holds more generally for a countably infinite group $G$ by a similar proof), we have $\Sigma(s)=\{c\}\cup\{gs\colon g\in G\}$. Consequently, a dynamic property is stable for $\sigma_{s}$ simply means that the NUCA $\sigma_{s}$ and the CA $\sigma_{c}$ satisfy the same property.
@@ 1.2 Main results
The situation for the surjunctivity of NUCA admits some complications. While injective and even reversible NUCA with finite memory may fail to be surjective (see [38, Example 14.3]), results in [38] show that reversible local perturbations of CA with finite memory over an amenable group or a residually finite group universe must be surjective. More generally, we obtain an extension (Theorem A) to cover all reversible, or equivalently stably injective, local perturbations of CA over surjunctive group universes. We also strengthen the conclusion by showing that such NUCA must be stably invertible (see Section 4).
@@
Combining with Theorem 1.1 and the result in [38, Theorem B] where we can replace the stable injectivity by the weaker injectivity condition whenever $G$ is an amenable group, we obtain the following general surjunctivity and invertibility result for NUCA which are local perturbations of CA.
@@
Let $M$ be a finite subset of a countable group $G$. Let $A$ be a finite alphabet and $S=A^{A^{M}}$. Let $s\in S^{G}$ be an asymptotically constant configuration. Then $\sigma_{s}$ is stably invertible in each of the following cases:
@@
Our next results concern a certain dual surjunctivity version of Gottschalk’s conjecture introduced by Capobianco, Kari, and Taati in [8] which states that every post-surjective CA over a group universe and a finite alphabet is also pre-injective. The authors settled in the same paper [8] the case of sofic group universes. See also [37] for some extensions.
@@
In particular, every sofic group is post-injunctive. By a similar technique as in the proof of Theorem A, we establish (see Section 5) the following extension of the above result of Capobianco, Kari, and Taati to cover the class of stably post-surjective local perturbations of CA over post-injunctive group universes.
@@
Note that residually finite groups are sofic and thus post-injunctive. It follows that [42, Theorem D.(ii)] is a consequence of our Theorem B when the group universe $G$ is residually finite.
Our main results also hold for global perturbations of CA with uniformly bounded singularity. A NUCA $\sigma_{s}\colon A^{G}\to A^{G}$ has uniformly bounded singularity if for every finite subset $E\subset G$ with $1_{G}\in E=E^{-1}$, we can find a finite subset $F\subset G$ containing $E$ such that the restriction $s|_{FE\setminus F}$ is constant. Note that in this paper, group operations are to be performed before boolean set operations. In particular, $FE\setminus F=(FE)\setminus F$. It is clear that $\sigma_{s}$ has uniformly bounded singularity if $s$ is asymptotically constant. When the universe $G$ is a residually finite group, our notion of uniformly bounded singularity is closely related but not equivalent to the notion of (periodically) bounded singularity for NUCA defined in [38, Definition 10.1]. For NUCA with uniformly bounded singularity, we obtain the following generalization of Theorem A (see Section 6)
@@
Let $M$ be a finite subset of a finitely generated surjunctive group $G$. Let $A$ be a finite alphabet and let $S=A^{A^{M}}$. Suppose that $\sigma_{s}\colon A^{G}\to A^{G}$ is stably injective for some $s\in S^{G}$ with uniformly bounded singularity. Then $\sigma_{s}$ is stably invertible.
Similarly, we can extend Theorem B to the context of NUCA with uniformly bounded singularity (see Section 7).
@@
Let $G$ be a finitely generated post-injunctive group and let $A$ be a finite alphabet. Let $M\subset G$ be finite and let $S=A^{A^{M}}$. Suppose that $\sigma_{s}\colon A^{G}~\to~A^{G}$ is stably post-surjective for some $s\in S^{G}$ with uniformly bounded singularity. Then $\sigma_{s}$ is stably invertible.
@@ 1.3 Perspectives
Our main results motivate the following natural questions which have an affirmative answer for surjunctive group universes and post-injunctive group universes respectively both in the setting of NUCA with finite memory and uniformly bounded singularity.
@@
Let $G$ be a group universe and let $A$ be a finite alphabet. Let $M\subset G$ be finite and let $S=A^{A^{M}}$. Suppose that $\sigma_{s}\colon A^{G}\to A^{G}$ is stably injective for some configuration $s\in S^{G}$. Is $\sigma_{s}$ stably invertible?
@@
Let $G$ be a group universe and let $A$ be a finite alphabet. Let $M\subset G$ be finite and let $S=A^{A^{M}}$. Suppose that $\sigma_{s}\colon A^{G}\to A^{G}$ is stably post-surjective for some configuration $s\in S^{G}$. Is $\sigma_{s}$ stably invertible?
Investigating the relation between these two questions, we can show (see Corollary 2) that Question 1 and Question 2 are in fact equivalent when restricted to the class of linear NUCA with finite memory. For every (possibly infinite) vector space $V$, note that $V^{G}$ is a vector space with component-wise operations and a NUCA $\tau\colon V^{G}\to V^{G}$ is said to be linear if it is also a linear map of vector spaces or equivalently, if its local transition maps are linear. Proposition 1 below provides a more precise relation between Question 1 and Question 2 than just their equivalence in the setting of linear NUCA with finite memory. For the proof of Proposition 1, we will need the following notion of dual NUCA introduced in [42].
Let $M\subset G$ be a finite subset and let $A$ be a vector space. Let $s\in S^{G}$ where $S=\mathcal{L}(A^{M},A)$ denotes the space of linear maps from $A^{M}$ to $A$. By linearity, we can define the linear endomorphism $s(g,m)\in\operatorname{End}(A)$ for all $m\in M$ and $g\in G$ by the formula
$s(g)(v)=\sum_{m\in M}s(g,m)v(m)\quad\text{ for all }v\in A^{M}.$ (1.1)
Setting $s(g,m)=0$ for all $m\in G\setminus M$, we can write for all $v\in A^{G}$ that
$s(g)(v)=\sum_{m\in G}s(g,m)v(m).$
Let $A^{*}$ be the dual space of $A$ and let $T=\mathcal{L}(A^{*M^{-1}},A^{*})$. We use the right superscript T to denote the transpose of linear maps. The dual configuration of local defining maps $s^{*}\in T^{G}$ is given by $s^{*}(g,m)\coloneqq s(gm,m^{-1})^{\mathsf{T}}$ for all $g,m\in G$. In particular, $s^{*}(g,m)=0$ for all $g\in G$ whenever $m^{-1}\notin M$ or equivalently when $m\notin M^{-1}$. Therefore, for all $u\in(A^{*})^{M^{-1}}$ and $g\in G$, we have
$s^{*}(g)(u)=\sum_{m\in M^{-1}}s^{*}(g,m)u(m)=\sum_{m\in M^{-1}}s(gm,m^{-1})^{\mathsf{T}}u(m).$ (1.2)
The linear NUCA $\sigma_{s}^{*}\coloneqq\sigma_{s^{*}}$ is called the dual linear NUCA of $\sigma_{s}$. We note that $s^{**}=s$ and $\sigma_{s}^{**}=\sigma_{s}$ for all $s\in S^{G}$ (see [42, Lemma 5.2]).
@@
Let $M$ be a finite subset of a group $G$ and let $A$ be a finite vector space alphabet. Let $S=\mathcal{L}(A^{M},A)$ and let $s\in S^{G}$. Then the following equivalences hold true:
@@
  4. (iv)
    $\sigma_{s}$ has uniformly bounded singularity if and only if so is $\sigma_{s^{*}}$.
  5. (v)
    $s$ is asymptotically constant if and only if so is $s^{*}$}.
Proof
The equivalences (i), (ii), and (iii) result directly from [42, Theorem A]. For (iv), let $E\subset G$ be a finite set such that $1_{G}\in E=E^{-1}$. Let $K\subset G$ be any finite subset large enough such that $1_{G}\in K=K^{-1}$ and $MEM^{-1}\subset K$. Suppose first that $\sigma_{s}$ has uniformly bounded singularity. Then there exists a finite subset $H\subset G$ such that $s|_{HK\setminus H}$ is constant. Let $F=HM$. We claim that $s^{*}|_{FE\setminus F}$ is also constant. Indeed, since $s|_{HK\setminus H}$ is constant and
$s(g)(v)=\sum_{m\in M}s(g,m)v(m)\quad\text{ for all }v\in A^{M},\,g\in G,$
there exists $t(m)\in\mathcal{L}(A^{M},A)$ for every $m\in M$ such that
$s(g,m)=t(m)\quad\text{ for all }g\in HK\setminus H,\,m\in M.$ (1.3)
Let $g\in FE\setminus F$ and $m\in M^{-1}$. In particular, $m^{-1}\in M$ and $g\notin F=HM$. Thus $g\notin Hm^{-1}$ and we deduce that $gm\notin H$. On the other hand, $m\in M^{-1}$ and $g\in FE=HME$ thus $gm\in HMEm\subset HMEM^{-1}$. As $MEM^{-1}\subset K$, it follows that $gm\in HK$. We conclude that $gm\in HK\setminus H$. We then infer from (1.3) that $s(gm,m^{-1})=t(m^{-1})$ for all $g\in FE\setminus F$ and $m\in M^{-1}$. Consequently, the formula (1.2) implies that for all $u\in(A^{*})^{M^{-1}}$, we have
$s^{*}(g)(u)=\sum_{m\in M^{-1}}s(gm,m^{-1})^{\mathsf{T}}u(m)=\sum_{m\in M^{-1}}t(m^{-1})^{\mathsf{T}}u(m)\quad\text{ for all }g\in FE\setminus F.$
Hence, $s^{*}|_{FE\setminus F}$ is constant. This proves that $\sigma_{s^{*}}$ has uniformly bounded singularity whenever $\sigma_{s}$ does. Since $(s^{*})^{*}=s$, the converse also holds and the proof of (iv) is complete.
For (v), suppose that $s$ is asymptotically constant, that is, $s|_{G\setminus H}$ is constant for some finite subset $H\subset G$. Then a similar (but simler) argument as above shows that $s^{*}|_{G\setminus HM}$ is constant. Since $HM$ is finite, this means that $s^{*}$ is also asymptotically constant whenever $s$ is asymptotically constant. As $(s^{*})^{*}=s$, the converse is true and we can conclude that $s$ is asymptotically constant if and only if so is $s^{*}$. ∎
Corollary 2
Let $G$ be a countable group and let $A$ be a finite vector space alphabet. Then Question 1 has an affirmative answer for all $\mathrm{NUCA}$ $\tau\colon A^{G}\to A^{G}$ in the class $C$ if and only if so does Question 2 where $C$ is one of the following:
  1. (a)
    $C=\{\text{linear }\mathrm{NUCA}\text{ with finite memory}\}$;
  2. (b)
[…]
  3. (c)
[…]
Proof
[…]
Linear NUCA with finite memory enjoy similar properties as linear CA or more generally group CA such as the shadowing property [3], [5], [29], [4], [30], [7], [36], [35], [12], [42]. The recent result [42, Theorem D.(ii)] states that every stably post-surjective linear NUCA with finite vector space alphabet over a residually finite group universe is invertible whenever it is a local perturbation of a linear CA. When specialized to the class of linear NUCA, Theorem B thus generalizes [42, Theorem D.(ii)] because every residually finite group is sofic and thus post-injunctive by Theorem 1.2. To this end, we postulate another seemingly interesting question.
@@
As a related problem, we can study analogues of our main results with different hypotheses on the configuration $s$ of local transition maps. For example, when the $G$-orbit of $s$ is finite, we obtain the following result (see Section 8).
Theorem E
Let $M$ be a finite subset of a countable group $G$. Let $A$ be a finite alphabet and let $S=A^{A^{M}}$. Suppose that $s\in S^{G}$ has finite $G$-orbit in $A^{G}$, i.e., $|Gs|=|\{gs\colon g\in G\}|<\infty$. The following hold:
  1. (i)
    If $G$ is surjunctive and $\sigma_{s}$ is injective, then $\sigma_{s}$ is invertible.
  2. (ii)
    If $G$ is post-injunctive and $\sigma_{s}$ is post-surjective, then $\sigma_{s}$ is invertible.
It is also interesting to extend computable properties of linear CA or more generally group CA [18], [19], [20] to the setting of linear or group NUCA.